On 1-1-2010 a machinery was purchased for 8000 rs it was should on 30-6-2013 for 6000rs. Depreciation is charged at 10% on original cost. Books are closed on 31st december every year prepare machinery account and depreciation account for the above perioud
1
Expert's answer
2020-03-02T08:43:45-0500

1. 8000*10%/100%=800 аnnual depreciation

8000/800=10 depreciation will be charged for 10 years from 30.06.2014 to 1.01. 2020

2.6000*10%/100%=600 annual despreciation

6000/600=10 depreciation will be charged for 10 years from 30.06.2013 to 30.06. 2023. Only in the half-year from 30.06.2013 to 01.01.2014 and 01.01.2023 to 30.06.2023 will be charged 300, since the incomplete year.